OpenStudy (anonymous):

how do i solve sin^2(x)>=0 ?

OpenStudy (anonymous):

so is it sin^(x)= 0?

OpenStudy (zzr0ck3r):

(anything)^2 >= 0 unless its imaginary

OpenStudy (anonymous):

prove \[\frac{\sin(2x)}{\sqrt{2-2\cos(2x)}}=\] and find any restrictions

OpenStudy (zzr0ck3r):

prove it?

OpenStudy (zzr0ck3r):

only resriction is 2-2cos(2x) > = 0

OpenStudy (anonymous):

= sin(x) my bad

OpenStudy (zzr0ck3r):

lol ok reask the question:)

OpenStudy (anonymous):

can i copy paste equation ??